Two trains set off at the same time from each end of a single straight railway line, which is 100km long. The first train travels at a constant speed of 55km/hour. the second train travels at a constant speed of 70km/hr. A very fast fly, Fred, travels at 100km/hr. Fred starts off in front of the first train and flys continuously back and forth between the two trains. How far does Fred fly before he is squashed between the two trains?

Answer by mananth(16949)

Train I speed = 55 km/h
Train II speed = 70 km/h
They are moving towards each other
so add up the speeds.
combined speed = 125 km/h
Distance = 100 km
Time taken by the trains to meet will be t= d/s
100/125 = 4/5 of an hour
During this time the fly is flying continuously.
speed = 100 km/h
distance traveled by Fred = speed * time
100 * 4/5 = 80 km
m.ananth@hotmail.ca